OpenStudy (anonymous):

Factor: x^2-13x+36 A. (x-3)(x-12) B. (x-4)(x-9) C. (x-6)

OpenStudy (anonymous):

9*4 = 36

OpenStudy (anonymous):

(x-4)(x-9) so B

OpenStudy (anonymous):

= x^2 - 4x - 9x + 36 = (x^2 - 4x) + (-9x + 36) = x(x-4)-9(x-4) = (x - 4) (x-9)